Find two consecutive odd intergers such that when the smaller is added to twice the larger the, the result is 24 more than the larger integer.
The formula I used was this x+x+2=24.

Answer by Paul(988)

Let the two consective odd integers be x and x+2
x+2(x+2)=24+(x+2)
x+2x+4=24+x+2
3x+4=26+x
2x=22
x=11

11+2=13

Hence, the two odd integers are 11 and 13.
Paul.